# VAMP-Net: An Interpretable Multi-Path Framework of Genomic Permutation-Invariant Set Attention and Quality-Aware 1D-CNN for MTB Drug Resistance

### v1

@@ 5 Experimentation and Discussion
This section presents the empirical validation of the Variant-Aware Multi-Path Network (VAMP-Net) and its core components, followed by a detailed discussion of the biological insights derived from our interpretability framework. We first detail the preparation of the dataset and the specific encoding utilized for the ChromPos_Ref>Alt variant tokens. Subsequent subsections systematically evaluate three critical aspects of the design of VAMP-Net: the selection of the optimal fusion mechanism, a comprehensive performance comparison, and a deep-dive interpretation of the predictions of the final model, including variant importance and epistatic interaction networks. The goal of this analysis is not only to demonstrate superior classification performance but to validate the translational utility of our dual-path approach for genetic resistance prediction.
@@ 5.1 Data Preprocessing
Prior to use, the data is subjected to multiple preparation procedures. After downloading the CRyPTIC_reuse_table_20240917.csv file from the official compendium website, we chose to focus on two drugs: Rifampicin (RIF), which is primarily used to treat tuberculosis by inhibiting RNA polymerase, and Rifabutin (RFB), a related rifamycin derivative with similar therapeutic applications.
First, we filter the dataset to include only high and medium quality samples, based on the {drug}_PHENOTYPE_QUALITY field. This results in 10,310 samples for RIF and 11,392 samples for RFB. For each selected sample, we download its corresponding VCF file, which contains variant information. Then we remove columns with missing data and retain only entries marked as PASS in the FILTER field, indicating reliable variant calls.
@@
We also retain additional descriptive features for each variant: [’GT’, ’DP’, ’DPF’, ’COV’, ’FRS’, ’GT_CONF’, ’GT_CONF_PERCENTILE’]. Numeric features are scaled using min–max normalization, while categorical features are encoded. The FRS field, representing the fraction of reads supporting the genotype call (#reads supporting GT / total reads), ranges between 0 and 1; Higher FRS values indicate stronger confidence in the genotype, whereas missing or low values typically correspond to low coverage or filtered calls, and hence we chose to replace its missing values with zeros. For each sample, these features are organised into an array of shape (number of variants × 8 features × 1), referred to as the Path-2 Input, which will be used later in a 1D-CNN model. Arrays are padded as needed to ensure uniform length across samples.
@@
Figure 4: Distribution of classes for the selected drugs.
Figure 4 shows the class distribution of the two drugs, with a difference of $\sim$17. 6% between the susceptible and resistant class for RIF, and $\sim$30. 4% for RFB, requiring us to employ a weighted cross-entropy loss to handle this class imbalance.
We will first use the RIF data in all subsequent experiments, navigating several design choices to determine the optimal approach, and subsequently apply the selected method to RFB for validation.
5.2 Variants Encoding
As mentioned in Section 5.1, variants are represented as Pos_REF>ALT, with the ALT for each allele separated by a comma. We compare a Static Encoding method , where each variant is given a unique ID to represent it, against training a BERT Tokenizer on all unique variants , and then using it to encode each sample’s variants into vectors before training.
The results showed that Static Encoding led to overfitting in the long run, with early stopping triggered at epoch 3. The training curve clearly showed signs of overfitting, with a $\sim$6.73% gap between training and validation accuracy.In contrast, using BERT proved effective in delivering more stable learning and improved validation performance. Although it also showed signs of overfitting, with a 4.83% difference between training and validation accuracy at the end of training, it maintained a smoother learning curve and reached early stopping at epoch 16.
The BERT-based encoding was more robust and critical for stable learning as shown in Figure 5. Consequently, we selected BERT for the encoding phase in all subsequent experiments.
Figure 5: Comparison of Static Encoding vs BERT Tokeniser metrics over epochs
5.3 The Choice of Fusion Mechanism
We tested three types of fusion strategies between Path-1 and Path-2, namely: Suppression, Amplification, and Suppression-Amplification (see Section 4.2.3). The training and validation curves are shown in Figure 6 and the test results are summarizsed in Table 1.
All three types achieved comparable performance in terms of validation and test accuracy. However, in general, the Amplification fusion yielded the best results and, therefore, it was selected as the preferred mechanism for the final configuration, and we will refer to this model as - Model A -.
Table 1: Comparison of fusion mechanisms. Best results in bold
@@ 5.3 The Choice of Fusion Mechanism
Figure 6: Training (top) and validation (bottom) performance curves for fusion methods
5.4 Models Experimentation
Starting from - Model A - , we evaluated a set of configurations to measure the robustness of the padding/masking strategy, data augmentation, and to assess the benefits of including Path-2 in the architecture. Moreover, we compared with a traditional approach, adopted in many studies, which involved representing the data as a matrix of shape (Number of VVariants) $\times$ (Number of Samples), where cell $i,j$ is 1 if Variant-$i$ is present in the sample, otherwise 0. This representation is typically trained with classifiers using either classical machine learning or deep learning techniques such as CNNs or MLP-based models. This hybrid of simple binary encoding and deep models has been widely surveyed in antimicrobial resistance (AMR) and clinical genomics literature and has been effective in practice (e.g., CNN and ML studies for tuberculosis and other pathogens) (sakagianni2024data). However, one limitation of this approach is the large resulting sparse matrix due to the high number of variants, i.e., the curse of dimensionality, requiring extensive computational resources. Some studies addressed this by pre-eliminating variants using statistical methods, or dimensionality reduction techniques like PCA (reshetnikov2025feature; pikalyova2024predicting; xu2025machine).
To select informative variants, we applied a Chi-squared test to assess the association between each variant and the target phenotype. Variants with $p$-value <0.001<0.001 were retained, resulting in 11,567 ($\sim$36.1% reduction), which were then used to train and evaluate both a multi-layer perceptron (MLP) and a convolutional neural network (CNN).
To systematically evaluate the contributions of the proposed components, we developed and tested several distinct model architectures:
@@ 5.4 Models Experimentation
  7. 7.
    Model-G: MLP for Presence/absence Variants data.
  8. 8.
    Model-H: CNN for Presence/Absence Variants data.
Table 2: Best hyperparameters of -Path-1 - for each proposed Multi-Path model applied on RIF Drug + Model A on RFB Drug
@@
Table 3: Best hyperparameters of - Path-2 - and the Classification head for each proposed Multi-Path model applied on RIF Drug + Model A on RFB drug
@@
Table 4: Best hyperparameters of MLP model aplied on RIF Drug

### v2

@@ 4 Experimentation and Discussion
This section presents the empirical validation of the Variant-Aware Multi-Path Network (VAMP-Net) and its core components, followed by a detailed discussion of the biological insights derived from our interpretability framework. We first detail the preparation of the dataset and the specific encoding utilized for the ChromPos_Ref>>Alt variant tokens. Subsequent subsections systematically evaluate three critical aspects of the design of VAMP-Net: the selection of the optimal fusion mechanism, a comprehensive performance comparison, and a deep-dive interpretation of the predictions of the final model, including variant importance and epistatic interaction networks. The goal of this analysis is not only to demonstrate superior classification performance but to validate the translational utility of our dual-path approach for genetic resistance prediction.
@@ 4.1 Data Preprocessing
Prior to use, the data is subjected to multiple preparation procedures. After downloading the CRyPTIC_reuse_table_20240917.csv file from the official compendium website, we focus on four drugs: Rifampicin (RIF), which is primarily used to treat tuberculosis by inhibiting RNA polymerase; Rifabutin (RFB), a related rifamycin derivative with similar therapeutic applications; Isoniazid (INH), a prodrug that targets the synthesis of mycolic acids, essential components of the mycobacterial cell wall; and Ethambutol (EMB), which acts by inhibiting the enzyme arabinosyltransferase, thereby disrupting the assembly of the mycobacterial cell wal.
First, we filter the dataset to include only high and medium quality samples, based on the {drug}_PHENOTYPE_QUALITY field. For each selected sample, we download its corresponding VCF file, which contains variant information. Then we remove columns with missing data and retain only entries marked as PASS in the FILTER field, indicating reliable variant calls.
@@
We also retain additional descriptive features for each variant: GT, DP, DPF, COV, FRS, GT_CONF, and GT_CONF_PERCENTILE. Numeric features are scaled using min–max normalization, while categorical features are encoded. The FRS field, representing the fraction of reads supporting the genotype call (#reads supporting GT / total reads), ranges between 0 and 1; higher FRS values indicate stronger confidence in the genotype, whereas missing or low values typically correspond to low coverage or filtered calls, and hence we chose to replace its missing values with zeros. For each sample, these features are organised into an array of shape (number of variants × 8 features × 1), referred to as the Path-2 Input, which will be used in a 1D-CNN model. Arrays are padded as needed to ensure uniform length across samples.
@@
Figure 5: Dataset Class Distribution across the selected anti-TB drugs. This figure illustrates the sample size ($n$) and the distribution between resistant (red) and susceptible (blue) classes for the four targeted drugs.
Figure 5 illustrates the sample counts and class distributions for each drug. We observe a class imbalance of approximately 17.6% between susceptible and resistant samples for RIF and 30.4% for RFB, while EMB exhibits the most significant disparity at 60.4%. To mitigate the effects of these imbalances during training, a weighted cross-entropy loss is employed
Initially, we utilize the RIF dataset to evaluate various architectural design choices and establish an optimal modeling pipeline. This finalized approach is subsequently applied to RFB, INH, and EMB to validate its generalizability across different drug resistance profiles. The train/validation/test split was (80/10/10), clarifying that this was performed at the isolate level to prevent data leakage. We also specify that stratification was applied based on the phenotypic resistance labels to ensure balanced representation across all sets.
4.2 Experimental Results
This section details a multi-stage evaluation of the VAMP-Net to quantify its predictive performance and justify its architectural design. We first establish a performance baseline by evaluating various feature encoding strategies and fusion strategies. A rigorous ablation study follows, isolating the specific contributions of the VCF-aware Path-2 branch to the model’s decision-making process. Finally, we report the performance of the best-performing VAMP-Net configuration across four primary anti-TB drugs (RIF, RFB, INH, and EMB), evaluating its robustness, calibration, and clinical utility on unseen isolates.
4.2.1 Evaluation of Feature Encoding Strategies
Following the preprocessing pipeline described in Section 4.1, genomic variants are represented as Pos_REFPos\_REF>ALTALT. We conducted a comparative evaluation between two distinct representational paradigms: (1) a Static Encoding approach, where each unique variant is mapped to a discrete identifier, and (2) a Contextualized Tokenization approach utilizing a BERT-based tokenizer trained on the corpus of unique variants to project genomic features into a continuous vector space.
Empirical results indicate that the Static Encoding method suffered from premature convergence and poor generalization. Early stopping was triggered at epoch 3, with the learning curves exhibiting a significant $\sim$6.73% generalization gap between training and validation accuracy. Conversely, the BERT-based representation demonstrated superior stability and robust feature extraction. While some degree of overfitting remained present—noted by a 4.83% accuracy disparity—the model exhibited a more sustained learning trajectory, converging at epoch 16 with a smoother loss profile. As illustrated in Figure 6, the BERT-based encoding facilitated more reliable optimization; consequently, it was selected as the foundational encoding strategy for all subsequent experiments.
Figure 6: Training and validation dynamics across tokenization strategies. Comparative analysis of BERT Tokenizer (top row) and Static Encoding (bottom row) across Accuracy, Loss, and ROC-AUC metrics. While both methods reach high training performance, the BERT-based approach demonstrates superior stability and generalization, as evidenced by the significantly narrower gap between training and validation curves in the Loss and ROC-AUC plots compared to the overfitting observed in Static Encoding.
4.2.2 Evaluation of Multimodal Late Fusion Strategies
To identify the optimal integration strategy between Path-1 and Path-2, we evaluated three distinct fusion paradigms: Suppression, Amplification, and a hybrid Suppression-Amplification mechanism. The optimization trajectories are illustrated in Figure 7, while a comprehensive performance summary across the test set is provided in Table 2.
Table 2: Comparative performance metrics across Gated Adaptive Fusion mechanisms. The table evaluates three functional modes of the adaptive fusion head: Suppression, Amplification, and a combined Suppression-Amplification strategy.
@@ 4.2.2 Evaluation of Multimodal Late Fusion Strategies
Figure 7: Impact of Gated Fusion strategies on model convergence during training. The trajectories compare three fusion variants: Amplification (orange), Suppression (yellow), and the combined Suppression-Amplification mechanism (green). While all methods converge to high ROC-AUC values (>0.95>0.95), the Amplification strategy exhibits the most stable optimization profile over 16 epochs, whereas the combined Suppression-Amplification approach shows greater variance, suggesting a trade-off between strict noise filtering and signal preservation.
While all three mechanisms demonstrated competitive performance with negligible variance in validation and testing metrics, the Amplification Fusion consistently yielded the highest predictive accuracy and exhibited superior generalization. Consequently, this strategy was adopted for the final model architecture, hereafter referred to as Model A.
4.2.3 Model Validation and Ablation Study
Building upon the selection of Model A as the architectural baseline, we conducted an extensive ablation study to evaluate the robustness of our padding/masking strategies, the efficacy of data augmentation, and the specific contributions of the dual-path integration. To systematically dissect the influence of each component, we evaluated six model configurations (Models A–F), as detailed below. The architectural variants were designed to isolate the effects of Attention Masking, Data Augmentation, and the Path-2 feature stream.
@@ 4.2.3 Model Validation and Ablation Study
Optimal hyperparameters were identified via Bayesian Optimization (Tables 3, 4), prioritizing a robust balance between the Area Under the Receiver Operating Characteristic (ROC AUC) curve and Balanced Accuracy on the validation set. This approach ensures that the selected configurations prioritize generalization over mere training set convergence.
Table 3: Comprehensive hyperparameter configurations for VAMP-Net. The table summarizes the optimized structural for Models A through F applied on RIF drug and model A apllied on RFB, INH and EMB drugs. This mapping includes the specific configurations for Path-1 Transformer (SAB).
@@
Model A (INH) 128 32 2
Model A (EMB) 128 32 2
Table 4: Comprehensive hyperparameter configurations for VAMP-Net. The table summarizes the optimized structural and learning parameters for Models A through F applied on RIF drug and model A apllied on RFB, INH and EMB drugs. This mapping includes the specific configurations for Path-2 (1D-CNN) and the Classification head.
@@
Model A (INH) 0.326 relu 1 3 0.00141
Model A (EMB) 0.326 relu 1 3 0.00141
The role of data augmentation, implemented via variant sequence shuffling, was scrutinized by comparing Model A and Model B. Shuffling leverages the Set Attention Block (SAB)’s inherent capacity for permutation invariance, ensuring that the model learns to identify critical genomic signatures independent of their relative position in the input sequence. As illustrated by the optimization trajectories in Figure 8, Model B exhibited superior training stability and a narrowed generalization gap, acting as an effective regularizer. On the test set, Model-A achieved the highest performance (ROC AUC: 0.969, Balanced Accuracy: 0.945). While Model-B’s test performance was marginally lower (ROC AUC: 0.964, Balanced Accuracy: 0.941), this minor trade-off in peak performance resulted in a substantially more robust and stable model. This suggests that data augmentation effectively improves generalization and stability with only a negligible decrease in maximum predictive power.
The results summarized in Table 5 provide critical insights into the contribution of each modular component within the VAMP-Net (VN) network. A comparative analysis yields several key findings, the full VAMP-Net (VN) configuration, which incorporates attention masking and localized feature extraction, achieves the highest overall performance across most metrics, specifically reaching an Accuracy of 0.956 and an F1-score of 0.960. This confirms that the synergistic integration of the dual-path architecture and precise masking provides the most robust predictive signal for Rifampicin resistance. Notably, the removal of the attention mask ("No mask, no shuffle") does not result in a catastrophic loss of performance, maintaining a high ROC-AUC of 0.973. This suggests that the SAB mechanism is inherently capable of identifying informative tokens and effectively disregarding padding, even without explicit masking. However, the slightly lower Balanced Accuracy (0.942) compared to the baseline (0.944) indicates that masking still provides a marginal benefit in stabilizing predictions for minority classes. The exclusion of Path-2 ("No Path-2") consistently resulted in a performance decrement, with Balanced Accuracy dropping to its lowest point (0.930). This validates the hypothesis that Path-B captures complementary genomic information, likely broader contextual or inter-variant relationships—that a single-path transformer architecture might overlook.
Figure 8: Validation of permutation invariance via variant sequence shuffling. Comparative analysis of model performance under data augmentation (top row) and standard training (bottom row). The top panels demonstrate that variant shuffling does not degrade performance; rather, it acts as a regularizer that tightens the generalization gap. The stability of ROC-AUC and Accuracy across both rows, despite the randomized ordering of input mutations, validates the permutation-invariant inductive bias of the Set Attention Block (SAB), ensuring the model captures biological topology regardless of variant positioning.
Table 5: Quantitative performance comparison of VAMP-Net architectural variants on the RIF test set. The table evaluates the influence of attention masking, data augmentation (shuffling), and dual-path integration across four key metrics.
Model variant Accuracy Bal. Acc. F1-score ROC-AUC
A: VAMP-Net (VN) 0.956 0.944 0.960 0.970
B: VN + shuffle 0.945 0.941 0.950 0.963
C: No mask, no shuffle 0.944 0.942 0.953 0.973
D: No mask + shuffle 0.945 0.943 0.953 0.969
E: No Path-2 + shuffle 0.940 0.930 0.940 0.961
F: No Path-2, no shuffle 0.940 0.935 0.941 0.960
4.2.4 VAMP-Net for anti-TB Drugs
The predictive performance of VAMP-Net across four primary anti-TB drugs is detailed in Table 6. To rigorously evaluate model stability, we report metrics across two dimensions: validation stability (top panel) and final unseen test set efficacy (bottom panel). Across three independent experimental seeds, VAMP-Net demonstrated high architectural stability. Standard deviations for all primary metrics remained consistently low, with ROC-AUC variance specifically constrained to <0.007<0.007. On the test set, INH achieved the highest overall performance, recording an accuracy of 0.945 $\pm$ 0.005 and a Matthews Correlation Coefficient (MCC) of 0.891 $\pm$ 0.010. RIF exhibited a similarly robust profile with an F1-score of 0.946 $\pm$ 0.007. For RFB, the model achieved its peak discriminative power with an ROC-AUC of 0.975 $\pm$ 0.001. A notable observation is the parity between Accuracy and Balanced Accuracy across most drugs (e.g., identical values of 0.945 for INH and a marginal 0.01 difference for EMB). Despite the inherent complexity of EMB resistance, VAMP-Net maintained the highest PR-AUC of the cohort at 0.977 $\pm$ 0.002. Collectively, the MCC values remained high across the spectrum (ranging from 0.614 to 0.891), indicating that the model maintains high-quality predictions across both resistant and susceptible classes.
Table 6: Statistical performance profile of VAMP-Net across first-line and second-line anti-TB drugs. The top panel provides validation stability metrics (Mean $\pm$ SD) across three independent experimental seeds to assess model convergence. The bottom panel details the final predictive performance on the unseen test set. We report Accuracy, Balanced Accuracy, F1-score, Matthews Correlation Coefficient (MCC), and Area Under the ROC and Precision-Recall Curves (ROC-AUC, PR-AUC).
Drug Accuracy Bal. Acc. F1 MCC ROC-AUC PR-AUC
RIF 0.937 $\pm$ 0.010 0.937 $\pm$ 0.013 0.946 $\pm$ 0.008 0.871 $\pm$ 0.022 0.967 $\pm$ 0.004 0.970 $\pm$ 0.005
RFB 0.925 $\pm$ 0.003 0.927 $\pm$ 0.002 0.924 $\pm$ 0.002 0.853 $\pm$ 0.003 0.971 $\pm$ 0.001 0.963 $\pm$ 0.002
EMB 0.839 $\pm$ 0.006 0.851 $\pm$ 0.007 0.892 $\pm$ 0.005 0.607 $\pm$ 0.007 0.915 $\pm$ 0.007 0.975 $\pm$ 0.002
INH 0.935 $\pm$ 0.007 0.935 $\pm$ 0.007 0.936 $\pm$ 0.007 0.870 $\pm$ 0.014 0.964 $\pm$ 0.000 0.953 $\pm$ 0.002
Test metrics:
Drug Accuracy Bal. Acc. F1 MCC ROC-AUC PR-AUC RIF 0.937 $\pm$ 0.009 0.935 $\pm$ 0.013 0.946 $\pm$ 0.007 0.870 $\pm$ 0.019 0.968 $\pm$ 0.006 0.970 $\pm$ 0.009 RFB 0.935 $\pm$ 0.008 0.937 $\pm$ 0.007 0.934 $\pm$ 0.006 0.874 $\pm$ 0.013 0.975 $\pm$ 0.001 0.960 $\pm$ 0.003 EMB 0.843 $\pm$ 0.014 0.853 $\pm$ 0.009 0.895 $\pm$ 0.010 0.614 $\pm$ 0.022 0.920 $\pm$ 0.004 0.977 $\pm$ 0.002 INH 0.945 $\pm$ 0.005 0.945 $\pm$ 0.005 0.945 $\pm$ 0.005 0.891 $\pm$ 0.010 0.970 $\pm$ 0.002 0.961 $\pm$ 0.004
Figure 9: Comparative test set performance across first-line and second-line anti-TB drugs. The bar chart illustrates the mean performance metrics across three independent experimental seeds for RIF, RFB, INH, and EMB. Error bars represent the standard deviation ($\pm$ SD), indicating high model stability across different initializations. Star markers ($\star$) denote the peak performance achieved for specific drug-metric pairs.
To evaluate the clinical readiness of VAMP-Net, we conducted a multi-dimensional analysis focusing on predictive accuracy, computational throughput, and stochastic stability across four major anti-TB drugs. As illustrated in Figure 9, VAMP-Net demonstrated high discriminative performance across the cohort. The model achieved peak ROC-AUC values exceeding $0.96$ for RIF, RFB, and INH, with INH specifically reaching a mean Accuracy and Balanced Accuracy of $0.945\pm 0.005$. While EMB presented the most challenging classification task, indicated by a lower Matthews Correlation Coefficient (MCC) of $0.614$, it maintained a superior Precision-Recall AUC ($0.977$), suggesting that the model remains highly effective at identifying true resistant isolates despite significant class imbalance.
The calibration of our model’s probabilistic outputs was assessed using the Brier Score (Figure 10 Left). VAMP-Net exhibited high calibration stability for INH ($\sim$0.05) and moderate calibration for RIF and RFB. The higher Brier score for EMB ($\sim$0.14) aligns with its lower MCC, further highlighting the biological complexity inherent in Ethambutol resistance pathways. Simultaneously, inference speed was evaluated to determine suitability for clinical deployment. VAMP-Net maintained a high-throughput processing rate of $\sim$490 samples per second across all drugs ( Figure 10 Right), ensuring near-instantaneous diagnostics even for large genomic surveillance datasets. To ensure the reported metrics were not artifacts of favorable initialization, we measured per-seed variability (Figure 11). The results show tight clustering around the mean for ROC-AUC, F1-score, and MCC across three independent seeds. The negligible variance in RIF and INH performance underscores the architectural robustness of our attention-based framework.
Figure 10: Computational efficiency and calibration across targeted anti-TB drugs. Left: Boxplots illustrate the distribution of Brier scores across three independent seeds, where a lower score indicates superior model calibration and probabilistic accuracy. Right: Throughput analysis measured in samples per second during inference. VAMP-Net maintains a consistent high-speed processing rate across all drugs, averaging $\sim$490 samples/s. Error bars represent the standard deviation ($\pm$ SD) across three experimental runs.
Figure 11: Per-seed variability of key performance metrics across three independent experimental runs. The scatter-boxplots illustrate the distribution of ROC-AUC (left), F1-score (center), and Matthews Correlation Coefficient (MCC, right) for RIF, RFB, INH, and EMB. Individual points represent specific stochastic seeds (Seed 42, Seed 123, and Seed 2024), while the diamond marker ($\diamond$) denotes the arithmetic mean. Error bars indicate the range of variability across runs. The high degree of clustering around the mean for RIF, RFB, and INH demonstrates the architectural stability of VAMP-Net. While EMB exhibits slightly higher dispersion, reflecting its greater biological and dataset complexity, the consistent performance across different initializations underscores the robustness of the Set Attention mechanism in capturing significant resistance signatures.
4.3 Comparative Analysis
4.3.1 Comparative Performance Against Baseline Architectures
We benchmarked our approach against the traditional binary encoding paradigm prevalent in antimicrobial resistance (AMR) and clinical genomics literature [36]. In this conventional framework, genomic data is represented as a high-dimensional sparse matrix $\mathbf{X}\in\{0,1\}^{N\times V}$, where $X_{i,j}$ denotes the presence or absence of variant $j$ in sample $i$. While effective in certain contexts, this representation suffers from the curse of dimensionality, necessitating significant computational resources or aggressive feature selection, such as PCA or statistical pruning, to manage the sparse feature space [34, 33, 48]. To ensure a fair comparison, we performed feature selection using a Chi-squared test ($\alpha=0.001$) to retain significantly associated variants, resulting in a reduced set of 11,567 features ($\sim$36.1% reduction) used to train VAMP-Net and competitive Multi-Layer Perceptron (MLP) and Convolutional Neural Network (CNN) baselines applied on RIF drug. The optimazed hyperparameter configurations for MLP and CNN are presented in Tables 7, and 8 respectively.
Table 7: Optimized hyperparameter configuration for the MLP baseline. The table outlines the best-performing structural parameters determined via empirical tuning on Rifampicin (RIF) resistance prediction.This specific configuration is referred to as Model G